OpenStudy (anonymous):

Can someone show me how to do this integral problem? 3x sqrt -x^2 + 1 dx

OpenStudy (anonymous):

Yeah Me

OpenStudy (akshay_budhkar):

whats ur ques? type it in equation no

OpenStudy (anonymous):

\[\int 3x \sqrt(-x^2 +1)dx\]

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

try \[u=1-x^2\]

OpenStudy (anonymous):

Well put \[t = -x^2 +1\]then differentiate

OpenStudy (akshay_budhkar):

lol

OpenStudy (anonymous):

to get \[\frac{dt}{dx} =-2x\] \[dx = \frac{dt}{-2x}\]

OpenStudy (anonymous):

Then just substitution

OpenStudy (anonymous):

\[ \frac{3}{-2} \int \sqrt t dt \]

OpenStudy (anonymous):

thats easy enough to integrate

OpenStudy (akshay_budhkar):

good one.